OpenStudy (anonymous):

Simplify: x^2 - 36 -------- x^2 - 12x + 36

OpenStudy (anonymous):

Choices: x - 6 ---- x - 3 x + 6 ---- x - 3 x - 3 ---- x + 3 x + 3 ----- x + 6

OpenStudy (anonymous):

hint: x^2 - 36 = x^2 - (6)^2 = (x+6)(x-6) x^2 - 12x + 36 = (x-6)(x-6)

OpenStudy (anonymous):

what xlegendx wrote is correct perhaps your wrote the wrong equation down or the wrong answers down? Is there a co-efficient in front of the x^2 for the denominator?